the equation whoose roots are sec2a & cosec2a
can be{where 2 is the power of sec &cosec}.

Grade:

1 Answers

Nehal Wani
14 years ago

Sum Of Roots
=(1/sin2a)+(1/cos2a)=(sin2a+cos2a)/(sinacosa)2=4/(sin2a)2=4cosec22a

Product Of Roots
=(1/sin2a)(1/cos2a)=1/(sinacosa)2=4/(sin2a)2=4cosec22a

X2-(Sum Of Roots)x+Product Of Roots=0

X2-(4cosec22a)x+4cosec22a=0